OpenStudy (anonymous):

There is available 50 liters of a 40% solution of acid and water. What volume of water must be added to reduce the concentration to 10% ?

OpenStudy (amistre64):

a + 50 = t <-- quantities a(1) + 50(.40) = t(.10) <-- concentrations

OpenStudy (anonymous):

150 liters?

OpenStudy (amistre64):

that and water has 0% acidity not 100% :)

OpenStudy (amistre64):

a + 50 = t ;*(-.10) a(0) + 50(.40) = t(.10) a(-.10)+ 50(-.10)= t(-.10) a (0) + 50(.40) = t(.10) ----------------------- a(0-.10) +50(.4-.1) = 0 a(-.1) +50(.3) = 0 a(-.1) +15 = 0 a = 15/.1 = 150

OpenStudy (anonymous):

Again, I just solved \[ x-0.9x = 15 \]

OpenStudy (amistre64):

pffft!! thats how they do it in kindergarten. At my age we do "real" math ;)

OpenStudy (anonymous):

between two what si the correct answer ?..

OpenStudy (amistre64):

our answers seem to converge on 150; but i havnt taken the limit of that yet

OpenStudy (anonymous):

lol @amistre64

OpenStudy (anonymous):

ok tnx ... :-)

OpenStudy (anonymous):

Glad to help :)

OpenStudy (anonymous):

40% * 50 + water = 10% ( 50 + L ) -> L = 200 - 50 = 150 L of water